# Broadening of the Divertor Heat Flux Profile in High Confinement Tokamak Fusion Plasmas with Edge Pedestals Limited by Turbulence in DIII-D

### v2

@@ Abstract
The XGC simulations of toroidal mode number N>5N\!>5\! modes show overall consistency with the fluctuation measurements described above. Fig. 8(a) shows $\delta n/n$ for the dominant toroidal mode number ($N$) fluctuations propagating in the electron direction, for WPQH-mode ($N=7$) and QH-Mode ($N=9$), indicating that XGC simulations find a relative increase for this toroidal mode similar to that measured in Figs. 4(a) in the steep gradient pedestal region. We have implemented synthetic BES in XGC to attempt validation of the XGC results against experimental observations at higher frequencies corresponding to $N>5$; preliminary results nearly match the total measured BES $\delta n/n$ profile for the QH case. The corresponding cross-phase between density and potential fluctuations $(\delta n,\delta\phi)$ from XGC is shown in Fig. 8(b). Noting that an “adiabatic” or Boltzmann electron density response would be in phase with the electrostatic potential, a strong non-adiabatic electron behavior (typical of the trapped electron response) is indicated by the non-zero cross-phase. This produces significant electron particle and thermal transport, typical of TEM turbulence. From XGC, the lab frame frequency of the QH feature is 330 kHz, while the poloidal wavenumber $k_{\theta}=0.63$ cm-1, consistent with the measurements in Fig. 4(c, d). In the plasma frame moving with the $E\times B$} velocity, the WPQH-Mode case shows both ion and electron modes spanning a range of wavenumbers, $k_{\theta}\sim 0.1-0.6$ cm-1 (Fig. 8(c)), while in contrast, the QH-Mode case shows mainly electron modes with $k_{\theta}\sim 0.4-0.7$ cm-1 (Fig. 8(d)).
@@
Finally, scaling arguments based on theory and simulation suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to naturally maintain stability to edge localized modes (ELMs) [ernst:2018, ernst:2022aps] so that Non-ELMing regimes naturally arise (see Appendix for details). Intrinsically Non-ELMing operating regimes such as Wide Pedestal Quiescent H-Mode and Quiescent H-Mode offer a high-performance solution without ELMs at relevant low pedestal top collisionalities, with the added benefit of broadened divertor heat flux profiles when the pedestal turbulent transport is sufficiently strong.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Award(s) DE-FC02-04ER54698, DE-SC0014264, DE-AC02-09CH11466, DE-SC0019004, DE-AC52-07NA27344, DE-NA0003525, DE-FG02-08ER54999, and DE-SC0019352. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
[…]
Appendix A Appendix
Eich Function. The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@ Appendix A Appendix
where $\bar{s}=s-s_{0}=(R_{\mathrm{mp}}-R_{\mathrm{LCFS}})f_{x}$ is the radial coordinate at the divertor, referring to the departure in midplane major radius from the LCFS, and $f_{x}=R_{\mathrm{div}}B_{\mathrm{pol}}^{\mathrm{div}}/R_{\mathrm{mp}}B_{\mathrm{pol}}^{\mathrm{mp}}\approx 5.3$ (for our cases) is the flux expansion effective area factor from midplane to divertor. The two parameters $(\lambda_{q},S)$ characterize the midplane exponential decay length and the Gaussian width, referred to the midplane, where $S$ represents competition between parallel and perpendicular heat transport between the X-point and strike point [eich:2013]. The Eich function is the convolution of an exponential decay in the scrape-off layer and a Gaussian diffusive width characterized by S. Note the IR data for these experiments has insufficient coverage to constrain S, which is obtained from Langmuir probe measurements.
@@
Figure A.1: Edge radial profiles for DIII-D #184833, comparing WPQH and QH-Mode phases. (a) Radial electric field $E_{r}$, (b) Hahm-Burrell $E\times B$} shearing rate, (c) electron density, (d) electron temperature, and (e) carbon ion temperature.
@@
Profile Fits. Pedestal profiles, shown in Fig. A.1, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge #184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
@@
Divertor Heat Flux Measurements. Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. A.3. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. A.2. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.
@@
Figure A.2: Comparison of $\lambda_{q}$ inferred from Langmuir probe (using $\gamma_{\mathrm{SH}}=7$) and IR thermography.
@@
To clarify the issue of possible heat flux sharing between the upper and lower divertors, While the overall shape is double-null characterized by two magnetic X-points, a -5 mm radial separation at the midplane between their corresponding flux surfaces sends most of the heat flux to the lower divertor [petrie:2006]. In all of the other, more recent, discharges used in this study the separation exceeds -10 mm, which well exceeds $\lambda_{q}$, so that heat flux sharing should not affect the inference of $\lambda_{q}$.
@@
Figure A.3: Measured parallel heat flux profiles from Infrared (IR) thermography (above, for cases where measurement is available), and Langmuir probes (below) for the discharges in Fig. 2(b). Solid lines show Eich fits with resulting parameters given in the insets. The dashed lines overlay the Eich IR fit on Langmuir probe profiles after adjusting to match $q_{\mathrm{BG}}$.
@@
Scaling Pedestal Turbulent Transport to Future Machines. Here we provide details on the scaling of pedestal turbulent transport with ion gyroradius, which varies inversely with magnetic field. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius, $\rho_{*}=\rho_{i}/a$, where $a$ is the plasma minor radius [kotschenreuther:1996]. In future machines, $\rho_{*}$ is expected to be approximately three times smaller than in present tokamaks. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ according to [zhang:1992, kotschenreuther:2017, hatch:2018]
@@
Here $\beta_{p}=8\pi p/B_{p}^{2}$ is the poloidal beta parameter, which we have included by invoking the EPED scaling for the pedestal width as limited by the onset of kinetic ballooning modes, $\Delta\sim\beta_{p}^{\alpha_{1}}$, where $\alpha_{1}\sim 0.5-0.75$ [snyder:2011, snyder:2012]. Assuming an H-Mode pedestal is formed, pedestal turbulent transport could be approximately an order of magnitude stronger in future high magnetic field machines as a result of the factor of three reduction in $\rho_{*}$, assuming the pedestal width follows EPED scaling (KBM onset) or a “softer” version of it with $\alpha_{1}<0.75$ due to other less virulent instabilities.

### v5

@@ Abstract
The XGC simulations of toroidal mode number N>5N\!>5\! modes show overall consistency with the fluctuation measurements described above. Figure 8(a) shows $\delta n/n$ for the dominant toroidal mode number ($N$) fluctuations propagating in the electron direction, for WPQH-mode ($N=7$) and QH-mode ($N=9$), indicating that XGC simulations find a relative increase for this toroidal mode similar to that measured in Figs. 4(a) in the steep gradient pedestal region. The corresponding cross-phase between density and potential fluctuations $(\delta n,\delta\phi)$ from XGC is shown in Fig. 8(b). Noting that an “adiabatic” or Boltzmann electron density response would be in phase with the electrostatic potential, a strong nonadiabatic electron behavior (typical of the trapped electron response) is indicated by the non-zero cross-phase. This produces significant electron particle and thermal transport, typical of TEM turbulence. From XGC, the lab frame frequency of the QH feature is 330 kHz, while the poloidal wave number $k_{\theta}=0.63$ cm-1, consistent with the measurements in Figs. 4(c), 4(d). In the plasma frame moving with the $E\times B$} velocity, the WPQH-mode case shows both ion and electron modes spanning a range of wave numbers, $k_{\theta}\sim 0.1-0.6$ cm-1 (Fig. 8(c)), while in contrast, the QH-mode case shows mainly electron modes with $k_{\theta}\sim 0.4-0.7$ cm-1 (Fig. 8(d)).
Initial collisionless estimates using uniformly distributed Maxwellian tracer electrons in the region $0.98<\psi_{N}<1.0$ suggest that turbulent magnetic fluctuations acting alone are less important than fluctuations in the $E\times B$} velocity in causing these particles to hit the divertor. However, we have discovered a strong synergy in which $E\times B$} fluctuations nonlinearly amplify electron transport by magnetic fluctuations. This new synergistic transport process will be explored in future work. Homoclinic tangles [evans:2004a] are observed in the XGC simulations near the X point, and are more pronounced in the QH-mode case, but subdominant.
@@
Finally, scaling arguments suggest that in future machines, pedestal turbulent transport may be sufficiently limiting to prevent ELMs [ernst:2018, ernst:2022aps], so that non-ELMing regimes naturally arise (see Appendix D for details). Intrinsically non-ELMing operating regimes, such as wide pedestal quiescent H-mode and quiescent H-mode, offer a high-performance solution at relevant low pedestal top collisionalities, featuring divertor heat flux profiles broadened by turbulence.
@@ Acknowledgements.
We thank Morgan Shafer (Oak Ridge National Laboratory) and Tom Osborne (General Atomics) for extensive comments and efforts during the DIII-D internal review process; as well as both referees for their detailed and useful constructive comments. We would like to acknowledge Keith Burrell (General Atomics), Florian Effenberg (Princeton Plasma Physics Laboratory), Qiming Hu (Princeton Plasma Physics Laboratory), Al Hyatt (General Atomics), Zeyu Li (General Atomics), Adam McClean (Princeton Plasma Physics Laboratory), Tomas Odstrčil (General Atomics), Theresa Wilks (Massachusetts Institute of Technology), Guanying Yu (Univ. California, Davis), and Lei Zeng (Univ. California Los Angeles) for contributions to the experiments featured. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-FC02-04ER54698, DE-SC0014264, DE-AC02-09CH11466, DE-SC0019004, DE-AC52-07NA27344, DE-NA0003525, DE-FG02-08ER54999, and DE-SC0019352. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. This research also used resources, via the INCITE program, of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Data included in figures and data required to reproduce simulation results are available from the author upon request. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
Appendix A: Eich Function.— The integral heat flux profile width $\lambda_{\mathrm{int}}$ can be related to the Eich fit parameters $(\lambda_{q},S)$ by $\lambda_{\mathrm{int}}=\int(q_{\|}(s)-q_{\mathrm{BG}})ds/(q_{\|0})\approx\lambda_{q}+1.64\,S$ [makowski:2012]. The Eich function is given by
@@
where $\bar{s}=s-s_{0}=(R_{\mathrm{mp}}-R_{\mathrm{LCFS}})f_{x}$ is the radial coordinate at the divertor, referring to the departure in midplane major radius from the LCFS, and $f_{x}=R_{\mathrm{div}}B_{\mathrm{pol}}^{\mathrm{div}}/R_{\mathrm{mp}}B_{\mathrm{pol}}^{\mathrm{mp}}\approx 5.3$ (for our cases) is the flux expansion effective area factor from midplane to divertor. The two parameters $(\lambda_{q},S)$ characterize the midplane exponential decay length and the Gaussian width, referred to the midplane, where $S$ represents competition between parallel and perpendicular heat transport between the X point and strike point [eich:2013]. The Eich function is the convolution of an exponential decay in the scrape-off layer and a Gaussian diffusive width characterized by S. Note the IR data for these experiments has insufficient coverage to constrain S, which is obtained from Langmuir probe measurements.
@@
Figure 9: Edge radial profiles for DIII-D No. 184833, comparing WPQH and QH-mode phases. (a) Radial electric field $E_{r}$, (b) Hahm-Burrell $E\times B$} shearing rate, (c) electron density, (d) electron temperature, and (e) carbon ion temperature.
@@
Appendix B: Profile Fits.— Pedestal profiles, shown in Fig. 9, are constrained by Langmuir probe measurements of separatrix electron temperature $T_{e}^{\mathrm{s}ep}$ and pedestal carbon and deuterium charge exchange measurements, which show constant carbon concentration across the pedestal. Equilibrium reconstructions of discharge No. 184833 in WPQH and QH phases, including the bootstrap current and realistic heating profiles from interpretative transport simulations (TRANSP) are used as input to the XGC code.
@@
Appendix C: Divertor Heat Flux Measurements.— Using the standard sheath heat transmission coefficient $\gamma_{\mathrm{SH}}=7$ [Stangeby:2000] to relate total parallel heat flux from electrons and ions to Langmuir probe measured saturation current ($J_{\mathrm{sat}}$) and electron temperature ($T_{e}$) via $q_{\|}^{\mathrm{tot}}=\gamma_{\mathrm{SH}}J_{\mathrm{sat}}T_{e}/e$, close agreement is obtained between Langmuir probe measured divertor heat flux profiles and IR thermography measured heat flux profiles, as shown in Fig. 11. The IR measurements are shown for corresponding cases where available. This agreement translates to agreement within statistical measurement uncertainty in the $\lambda_{q}$ values inferred from Langmuir probes and IR, as shown in Fig. 10. Importantly, the $\lambda_{q}$ inferred from Langmuir probes is independent of the assumed $\gamma_{\mathrm{SH}}$. Because Langmuir probes measure mainly electrons, this result can be viewed as evidence that the electron contribution to the divertor heat flux underlies the broadening of $\lambda_{q}$ measured by IR.
@@
Figure 10: Comparison of $\lambda_{q}$ inferred from Langmuir probe (using $\gamma_{\mathrm{SH}}=7$) and IR thermography.
@@
To clarify the issue of possible heat flux sharing between the upper and lower divertors, While the overall shape is double-null characterized by two magnetic X points, for No. 184833 a -5 mm radial separation at the midplane between their corresponding flux surfaces sends most of the heat flux to the lower divertor [petrie:2006]. In all of the other more recent discharges used in this study, the separation exceeds -10 mm, which well exceeds $\lambda_{q}$, so that heat flux sharing should not affect the inference of $\lambda_{q}$.
@@
Figure 11: Measured parallel heat flux profiles from IR thermography (above, for cases where measurement is available), and Langmuir probes (below) for the discharges in Fig. 2(b). Solid lines show Eich fits with resulting paraminfraredeters given in the insets. The dashed lines overlay the Eich IR fit on Langmuir probe profiles after adjusting to match $q_{\mathrm{BG}}$.
@@
Appendix D: Scaling Pedestal Turbulent Transport to Future Machines.— Here we provide details on the scaling of pedestal turbulent transport with ion gyroradius, which varies inversely with magnetic field. Turbulent transport reduction by sheared flows is predicted to scale with the normalized ion gyroradius, $\rho_{*}=\rho_{i}/a$, where $a$ is the plasma minor radius [kotschenreuther:1996]. In future machines, $\rho_{*}$ is expected to be approximately 3 times smaller than in present tokamaks. The radial electric field $E_{r}$ in the edge pedestal is $Zen_{i}E_{r}\sim dp_{i}/dr\sim p_{i}/\Delta$, where $Zen_{i}$ is the ion charge density, $p_{i}=n_{i}T_{i}$ is the ion pressure with $T_{i}$ the ion temperature, $r$ is the minor radius, and $\Delta$ is the gradient scale length The $E\times B$} shear rate for turbulence is then $\gamma_{E}\sim B^{-1}dE_{r}/dr\sim E_{r}/(B\Delta)$ where $B$ is the magnetic field. Estimating the growth rate for ion scale drift-type instabilities as $\gamma_{\mathrm{lin}}\sim v_{\mathrm{thi}}/\Delta$, where $v_{\mathrm{thi}}=(2T_{i}/m_{i})^{1/2}$ is the ion thermal speed, the reduction of turbulent transport by $E\times B$} shear will depend on a parameter which scales as $\gamma_{E}/\gamma_{\mathrm{lin}}\sim\rho_{i}/\Delta\sim(a/\Delta)\rho_{*}$. The ratio of pedestal turbulent heat flux $Q$ to the gyroBohm flux $Q_{\mathrm{gB}}$ is predicted to increase asymptotically as a result of the decreased $E\times B$} shear at low $\rho_{*}$ according to [zhang:1992, kotschenreuther:2017, hatch:2018]
@@
Here $\beta_{p}=8\pi p/B_{p}^{2}$ is the poloidal beta parameter, which we have included by invoking the EPED scaling for the pedestal width as limited by the onset of kinetic ballooning modes, $\Delta\sim\beta_{p}^{\alpha_{1}}$, where $\alpha_{1}\sim 0.5-0.75$ [snyder:2011, snyder:2012]. Assuming an H-mode pedestal is formed, pedestal turbulent transport could be approximately an order of magnitude stronger in future high magnetic field machines as a result of the factor of three reduction in $\rho_{*}$, assuming the pedestal width follows EPED scaling (KBM onset) or a “softer” version of it with $\alpha_{1}<0.75$ due to other less virulent instabilities [li:2024].